OpenStudy (anonymous):

A recipe for molding dough calls for flour and salt in a ratio of 2 to 3. If you use 6 cups of salt, how much flour should you use?

OpenStudy (anonymous):

4 cups

OpenStudy (anonymous):

2/3=x/6 12=3x x=4

OpenStudy (anonymous):

4 cups

OpenStudy (anonymous):

\[\frac{\text{flour}}{\text{salt}}=\frac{2}{3} \]Replace salt with it's value 6, and solve for flour